# The Complexity of Minimum-Envy House Allocation Over Graphs

### v2

@@ 1.1 Contributions
Next, we show that even when the preference lists of each agent is at most $2$, the problem becomes $\mathsf{NP}$-hard. It is to be noted that Madathil et. al [15] proved NP-hardness for House Allocation (Optimal House Allocation of Agent Network on agent graphs which are complete graphs) when the preference list of each agent is at most $2$. Our NP-hardness result for Optimal House Allocation of Agent Network holds even when the agent graph is a $3$-regular graph. The same NP-hardness result shows that Optimal House Allocation of Agent Network, and consequently Optimally Happy House Allocation of Agent Network, is $\mathsf{NP}$-hard even when the number $|\mathcal{A}|=n$ of agents is equal to the number $|\mathcal{H}|=m$ of houses. This is in stark contrast to the House Allocation problem, which can be solved in polynomial time when $n=m$. These proofs appear in detail as Theorem 3.5 and Theorem 3.7, in Section 3.
@@ 1.2 Related Work
The House Allocation problem and its various extensions have been extensively studied across economics and computer science [2, 3, 5, 7, 10, 12], particularly in contexts that emphasize fairness and efficiency. We discuss the most relevant prior works to our setting.
@@
Abdulkadiroğlu and Sönmez [1] introduce the House Allocation with Existing Tenants model, which captures scenarios where some agents already occupy houses and may choose to keep them or participate in reallocation. They propose the Top Trading Cycles (TTC) mechanism, which guarantees allocations that are Pareto efficient, individually rational, and strategy-proof. Their model generalizes the classical house allocation setting by incorporating endowments and offering mechanisms that incentivize participation without loss. Focusing on the classical House Allocation problem under ordinal preferences, Gan, Suksompong, and Voudouris [6] develop a polynomial-time algorithm to determine whether an envy-free allocation exists when each agent ranks houses. They explore conditions under which envy-free allocations can or cannot be guaranteed and highlight cases where approximations or partial allocations become necessary. Their work contributes foundational results on the computational aspects of envy-freeness under structured preference inputs and motivates the need for relaxed fairness notions studied in later works.
@@
Hosseini et al. [11, 9] introduce the Graphical House Allocation model, wherein agents are vertices in a graph, and envy is permitted only along the graph’s edges. Assuming identical valuation functions across agents, they focus on minimizing the total envy, measured as the sum of envy across edges. They show that the problem generalizes the classical linear arrangement problem and establish hardness results even for simple graphs like disjoint paths and stars. To address tractability, they develop fixed-parameter algorithms based on graph properties like separability and provide polynomial-time solutions for special families such as paths and cycles. Choo, Ling, Suksompong, Teh, and Zhang [4] address the challenge of achieving envy-freeness in settings where such allocations may not naturally exist. They propose the use of subsidies—monetary compensations that accompany house allocations—to compute envy-free outcomes while minimizing the total subsidy required. They show that the problem is NP-hard in general but solvable in polynomial time when the number of houses differs from the number of agents by a constant, or when agents have identical utilities.
@@
Madathil, Misra, and Sethia. [15] provide a comprehensive study on minimizing envy in House Allocation, introducing three formal measures of envy—number of envious agents, maximum envy, and total envy. They define and analyze the corresponding computational problems (Optimal House Allocation, Egalitarian House Allocation, and Utilitarian House Allocation), establish strong hardness results even under restricted settings like binary preferences and bounded degrees, and design polynomial-time algorithms for structured instances such as extremal preferences. They also present fixed-parameter tractable algorithms based on agent and house types, formulate ILP models with empirical evaluations, and study the trade-off between fairness and welfare through the Price of Fairness, offering tight bounds and identifying cases where fairness incurs no welfare loss.
@@ 2 Preliminaries
Optimal House Allocation of Agent Network Input: A set $\mathcal{A}$ of $n$ agents, a graph $\mathcal{G}_{\mathcal{A}}$ on $\mathcal{A}$, a set $\mathcal{H}$ of $m$ houses, and agent preferences $(\mathcal{P}_{a})_{a\in\mathcal{A}}$. Question: Compute an allocation $\phi$ that minimizes |$\left|\{a\in\mathcal{A}\mid\mathcal{E}_{\mathcal{A}}^{\phi}(a)\neq\emptyset\}\right|$.
Optimally Happy House Allocation of Agent Network Input: A set $\mathcal{A}$ of $n$ agents, a graph $\mathcal{G}_{\mathcal{A}}$ on $\mathcal{A}$, a set $\mathcal{H}$ of $m$ houses, and agent preferences $(\mathcal{P}_{a})_{a\in\mathcal{A}}$. Question: Compute an allocation $\phi$ that minimizes |$\left|\{a\in\mathcal{A}\mid\mathcal{E}_{\mathcal{A}}^{\phi}(a)\neq\emptyset\}\right|$, and among all such minimum-envy allocations, maximizes |$\left|\{a\in\mathcal{A}\mid\phi(a)\in\mathcal{P}_{a}\}\right|$.
@@ 3 Introducing an Agent Graph: Contrast with the Classical Setting
Notice that the House Allocation problem studied by Madathil et. al [15] is a special case of Optimal House Allocation of Agent Network where the agent graph is a complete graph. In this section, we provide results for Optimal House Allocation of Agent Network that depict that this has a much more interesting complexity landscape in comparison to the House Allocation problem.
We first look into a tractable variant of House Allocation that remains tractable for Optimal House Allocation of Agent Network. Madathil et. al [15] show that when each agent prefers exactly one house, Optimal House Allocation of Agent Network on complete graphs is polynomial-time solvable. We start by extending this result to solving Optimal House Allocation of Agent Network on arbitrary graphs in polynomial time, when each agent prefers exactly one house i.e. $d=1$.
@@
4:$\mathcal{G}_{\mathcal{F}}\leftarrow$ complete bipartite graph with partitions $\mathcal{A}$, $\mathcal{H}$, and weights $w$
@@
Madathil et. al. [15] show that Optimal House Allocation of Agent Network for cliques is hard even when each agent prefers at most two houses, i.e. $d=2$. We show that even for the class of bipartite graphs, Optimal House Allocation of Agent Network remains hard even when each agent prefers at most two houses, i.e. $d=2$. This involves a reduction from the $\mathsf{NP}$-hard problem of CLIQUE, which is a nontrivial extension of the work by Madathil et. al. [15].
@@
(a) Instance of CLIQUE for $3$-regular graphs, $k=3$
@@
    Let =$\mathcal{H}=\{h_{v}\mid v\in V\}\cup\{h^{*}_{j}\mid j\in[\delta N+M-k]\}$. Therefore, $|\mathcal{H}|=N+\delta N+M-k=|\mathcal{A}|+(N-k)$.
@@
[…]
@@
We now look into agents which are not assigned dummy vertices by $\phi$. There are $|\mathcal{A}|-(\delta N+M-k)=(\delta N+M)-(\delta N+M-k)=k$ many such agents, all of the form $a^{1}_{v}$, $v\in V$. Let $a^{1}_{v_{1}},a^{1}_{v_{2}},\ldots,a^{1}_{v_{k}}$ be the agents which are not assigned dummy houses. For any $e\in E$, $e$ is incident on least one vertex in $v_{1},v_{2},\ldots,v_{k}$, if and only if the agent $a_{e}$ is envious. Therefore, the number of envious agents is precisely the number of edges which are incident on the vertices $v_{1},v_{2},\ldots,v_{k}$. Let $\gamma$ be the number of edges with both endpoints in $\{v_{1},v_{2},\ldots,v_{k}\}$. The number of edges incident on $\{v_{1},v_{2},\ldots,v_{k}\}$ is therefore equal to $k\delta-\gamma$. By assumption, the number of envious agents is at most $k\delta-\binom{k}{2}$. Thus $\gamma\geqslant\binom{k}{2}$, implying $v_{1},v_{2},\ldots,v_{k}$ is a clique.
@@
Moreover, Madathil et. al [15] provide polynomial-time algorithms for the following cases: (i) when $|\mathcal{A}|=|\mathcal{H}|$, (ii) when preferences $(\mathcal{P}_{a})_{a\in\mathcal{A}}$ follow an extremal structure. Preferences follow an extremal structure if there is an ordering of houses in =$\mathcal{H}=\{h_{1},h_{2},\ldots,h_{m}\}$ such that for every agent $a\in\mathcal{A}$, they prefer either some suffix of the houses, $\mathcal{P}_{a}=\{h_{i},h_{i+1},\ldots,h_{m}\}$, or some prefix of the houses $\mathcal{P}_{a}=\{h_{1},h_{2},\ldots,h_{i}\}$, for some $i\in[m]$.
@@
We provide a reduction from the $1/2$-Vertex Separator problem [16] on 3-regular graphs to Optimal House Allocation of Agent Network. The problem takes input a graph $G(V,E)$ and an integer $k$ and asks if there is a subset $S$ of $V$ that $|S|\leqslant k$ such that $V\setminus S$ can be partitioned into two equal-sized subsets $S_{1}$, $S_{2}$ such that no edge connects a vertex in $S_{1}$ with a vertex in $S_{2}$. This problem is shown to be $\mathsf{NP}$-complete even for 3-regular graphs by Müller and Wagner [16].
@@
    Let =$\mathcal{H}^{*}=\{h_{1},h_{2},\ldots,h_{t}\}$, where $t=\frac{n}{2}-\left\lfloor\frac{k}{2}\right\rfloor$. $\mathcal{P}_{a}=\mathcal{H}^{*}$ for all $a\in\mathcal{A}$.
@@
Now, say $k$ is odd. and let $(G,k)$ be a Yes-instance. Let $k=2\lambda+1$ for $\lambda\in\mathbb{Z}$. Note that since $G$ is 3-regular, $n=|V|$ must be even. Let $S$ be any separator that separates $V\setminus S$ into equal sized, disjoint, and non-adjacent parts $S_{1}$, $S_{2}$ and satisfying $|S|\leqslant k$. Therefore, $|S_{1}|=|S_{2}|$. Hence $|V|-|S|=|S_{1}|+|S_{2}|$, or $|S|=n-2|S_{1}|$, an even integer: say $|S|=2\beta$, for some integer $\beta$. Therefore, $0.52\beta\leqslant 2\lambda+1\implies\beta\leqslant\lambda+$ 0.5. Since $\beta,\lambda$ are integers, we must have $\beta\leqslant\lambda$; implying $|S|=2\beta\leqslant 2\lambda=2\left\lfloor\frac{k}{2}\right\rfloor$. Hence $(G,2\left\lfloor\frac{k}{2}\right\rfloor)$ is also a Yes-instance.
@@
[…]
@@ 4.1 Single Exponential Exact Algorithms
    If $a\notin C$, then $\phi(a)\notin\mathcal{P}_{a}$. Thus $\mathcal{F}_{a}$ cannot contain any house in $\mathcal{P}_{a}$. While, for $a\in C$, $\mathcal{F}_{a}$ can only contain houses in $\mathcal{P}_{a}$.
@@
1:$\xi\leftarrow n$triangleright In the worst case, all agents are envious
@@
23:         $E_{\mathcal{F}}\leftarrow\{\{a,h\}\mid a\in\mathcal{A},h\in\mathcal{F}_{a}$}; $\mathcal{G}_{\mathcal{F}}\leftarrow(\mathcal{H}\cup\mathcal{A},E_{\mathcal{F}})$triangleright Feasibility bipartite graph

### v3

# Minimizing Envy and Maximizing Happiness in Graphical House Allocation

@@ 1.1 Contributions
Next, we show that even when the preference lists of each agent is at most $2$, the problem becomes $\mathsf{NP}$-hard. It is to be noted that Madathil et. al [madathil2024cost] proved NP-hardness for House Allocation (Optimal House Allocation of Agent Network on agent graphs which are complete graphs) when the preference list of each agent is at most $2$. Our NP-hardness result for Optimal House Allocation of Agent Network holds even when the agent graph is a $3$-regular graph. The same NP-hardness result shows that Optimal House Allocation of Agent Network, and consequently Optimally Happy House Allocation of Agent Network, is $\mathsf{NP}$-hard even when the number $|\mathcal{A}|=n$ of agents is equal to the number $|\mathcal{H}|=m$ of houses. This is in stark contrast to the House Allocation problem, which can be solved in polynomial time when $n=m$. These proofs appear in detail as Theorem 3.5 and Theorem 3.7, in Section 3.
@@ 1.2 Related Work
The House Allocation problem and its various extensions have been extensively studied across economics and computer science [abraham2004pareto, beynier2019local, dai2024weighted, gerards2025envy, hosseini2023tight, kamiyama2021complexity], particularly in contexts that emphasize fairness and efficiency. We discuss the most relevant prior works to our setting.
@@
Abdulkadiroğlu and Sönmez [abdulkadirouglu1999house] introduce the House Allocation with Existing Tenants model, which captures scenarios where some agents already occupy houses and may choose to keep them or participate in reallocation. They propose the Top Trading Cycles (TTC) mechanism, which guarantees allocations that are Pareto efficient, individually rational, and strategy-proof. Their model generalizes the classical house allocation setting by incorporating endowments and offering mechanisms that incentivize participation without loss. Focusing on the classical House Allocation problem under ordinal preferences, Gan, Suksompong, and Voudouris [gan2019envy] develop a polynomial-time algorithm to determine whether an envy-free allocation exists when each agent ranks houses. They explore conditions under which envy-free allocations can or cannot be guaranteed and highlight cases where approximations or partial allocations become necessary. Their work contributes foundational results on the computational aspects of envy-freeness under structured preference inputs and motivates the need for relaxed fairness notions studied in later works.
@@
Hosseini et al. [hosseini2023graphical, hosseini2024graphical] introduce the Graphical House Allocation model, wherein agents are vertices in a graph, and envy is permitted only along the graph’s edges. Assuming identical valuation functions across agents, they focus on minimizing the total envy, measured as the sum of envy across edges. They show that the problem generalizes the classical linear arrangement problem and establish hardness results even for simple graphs like disjoint paths and stars. To address tractability, they develop fixed-parameter algorithms based on graph properties like separability and provide polynomial-time solutions for special families such as paths and cycles. Choo, Ling, Suksompong, Teh, and Zhang [choo2024envy] address the challenge of achieving envy-freeness in settings where such allocations may not naturally exist. They propose the use of subsidies—monetary compensations that accompany house allocations—to compute envy-free outcomes while minimizing the total subsidy required. They show that the problem is NP-hard in general but solvable in polynomial time when the number of houses differs from the number of agents by a constant, or when agents have identical utilities.
@@
Madathil, Misra, and Sethia. [madathil2024cost] provide a comprehensive study on minimizing envy in House Allocation, introducing three formal measures of envy—number of envious agents, maximum envy, and total envy. They define and analyze the corresponding computational problems (Optimal House Allocation, Egalitarian House Allocation, and Utilitarian House Allocation), establish strong hardness results even under restricted settings like binary preferences and bounded degrees, and design polynomial-time algorithms for structured instances such as extremal preferences. They also present fixed-parameter tractable algorithms based on agent and house types, formulate ILP models with empirical evaluations, and study the trade-off between fairness and welfare through the Price of Fairness, offering tight bounds and identifying cases where fairness incurs no welfare loss.
@@ 2 Preliminaries
Optimal House Allocation of Agent Network Input: A set $\mathcal{A}$ of $n$ agents, a graph $\mathcal{G}_{\mathcal{A}}$ on $\mathcal{A}$, a set $\mathcal{H}$ of $m$ houses, and agent preferences $(\mathcal{P}_{a})_{a\in\mathcal{A}}$. Question: Compute an allocation $\phi$ that minimizes $\left|\{a\in\mathcal{A}\mid\mathcal{E}_{\mathcal{A}}^{\phi}(a)\neq\emptyset\}\right|$.
Optimally Happy House Allocation of Agent Network Input: A set $\mathcal{A}$ of $n$ agents, a graph $\mathcal{G}_{\mathcal{A}}$ on $\mathcal{A}$, a set $\mathcal{H}$ of $m$ houses, and agent preferences $(\mathcal{P}_{a})_{a\in\mathcal{A}}$. Question: Compute an allocation $\phi$ that minimizes $\left|\{a\in\mathcal{A}\mid\mathcal{E}_{\mathcal{A}}^{\phi}(a)\neq\emptyset\}\right|$, and among all such minimum-envy allocations, maximizes $\left|\{a\in\mathcal{A}\mid\phi(a)\in\mathcal{P}_{a}\}\right|$.
@@ 3 Introducing an Agent Graph: Contrast with the Classical Setting
Notice that the House Allocation problem studied by Madathil et. al [madathil2024cost] is a special case of Optimal House Allocation of Agent Network where the agent graph is a complete graph. In this section, we provide results for Optimal House Allocation of Agent Network that depict that this has a much more interesting complexity landscape in comparison to the House Allocation problem.
We first look into a tractable variant of House Allocation that remains tractable for Optimal House Allocation of Agent Network. Madathil et. al [madathil2024cost] show that when each agent prefers exactly one house, Optimal House Allocation of Agent Network on complete graphs is polynomial-time solvable. We start by extending this result to solving Optimal House Allocation of Agent Network on arbitrary graphs in polynomial time, when each agent prefers exactly one house i.e. $d=1$.
@@
[…]
@@
Madathil et. al. [madathil2024cost] show that Optimal House Allocation of Agent Network for cliques is hard even when each agent prefers at most two houses, i.e. $d=2$. We show that even for the class of bipartite graphs, Optimal House Allocation of Agent Network remains hard even when each agent prefers at most two houses, i.e. $d=2$. This involves a reduction from the $\mathsf{NP}$-hard problem of CLIQUE, which is a nontrivial extension of the work by Madathil et. al. [madathil2024cost].
@@
(a) Instance of CLIQUE for 33-regular graphs, $k=3$
@@
    Let $\mathcal{H}=\{h_{v}\mid v\in V\}\cup\{h^{*}_{j}\mid j\in[\delta N+M-k]\}$. Therefore, $|\mathcal{H}|=N+\delta N+M-k=|\mathcal{A}|+(N-k)$.
@@
Forward direction. Let $S\subseteq V$ be a clique in $G$ of size $|S|=k$. Consider the allocation $\phi$ defined as follows:
@@
We now look into agents which are not assigned dummy vertices by $\phi$. There are $|\mathcal{A}|-(\delta N+M-k)=(\delta N+M)-(\delta N+M-k)=k$ many such agents, all of the form $a^{1}_{v}$, $v\in V$. Let $a^{1}_{v_{1}},a^{1}_{v_{2}},\ldots,a^{1}_{v_{k}}$ be the agents which are not assigned dummy houses. For any $\in E, ee$ is incident on least one vertex in $v_{1},v_{2},\ldots,v_{k}$, if and only if the agent $a_{e}$ is envious. Therefore, the number of envious agents is precisely the number of edges which are incident on the vertices $v_{1},v_{2},\ldots,v_{k}$. Let $\gamma$ be the number of edges with both endpoints in $\{v_{1},v_{2},\ldots,v_{k}\}$. The number of edges incident on $\{v_{1},v_{2},\ldots,v_{k}\}$ is therefore equal to $k\delta-\gamma$. By assumption, the number of envious agents is at most $k\delta-\binom{k}{2}$. Thus $\gamma\geqslant\binom{k}{2}$, implying $v_{1},v_{2},\ldots,v_{k}$ is a clique.
@@
Moreover, Madathil et. al [madathil2024cost] provide polynomial-time algorithms for the following cases: (i) when $|\mathcal{A}|=|\mathcal{H}|$, (ii) when preferences $(\mathcal{P}_{a})_{a\in\mathcal{A}}$ follow an extremal structure. Preferences follow an extremal structure if there is an ordering of houses in $\mathcal{H}=\{h_{1},h_{2},\ldots,h_{m}\}$ such that for every agent $a\in\mathcal{A}$, they prefer either some suffix of the houses, $\mathcal{P}_{a}=\{h_{i},h_{i+1},\ldots,h_{m}\}$, or some prefix of the houses $\mathcal{P}_{a}=\{h_{1},h_{2},\ldots,h_{i}\}$, for some $i\in[m]$.
@@
We provide a reduction from the $1/2$-Vertex Separator problem [muller1991alpha] on 3-regular graphs to Optimal House Allocation of Agent Network. The problem takes input a graph $G(V,E)$ and an integer $k$ and asks if there is a subset $S$ of $V$ that $|S|\leqslant k$ such that $V\setminus S$ can be partitioned into two equal-sized subsets $S_{1}$, $S_{2}$ such that no edge connects a vertex in $S_{1}$ with a vertex in $S_{2}$. This problem is shown to be $\mathsf{NP}$-complete even for 3-regular graphs by Müller and Wagner [muller1991alpha].
@@
    Let $\mathcal{H}^{*}=\{h_{1},h_{2},\ldots,h_{t}\}$, where $t=\frac{n}{2}-\left\lfloor\frac{k}{2}\right\rfloor$. $\mathcal{P}_{a}=\mathcal{H}^{*}$ for all $a\in\mathcal{A}$.
@@
Now, say $k$ is odd. and let $(G,k)$ be a Yes-instance. Let $k=2\lambda+1$ for $\lambda\in\mathbb{Z}$. Note that since $G$ is 3-regular, $n=|V|$ must be even. Let $S$ be any separator that separates $V\setminus S$ into equal sized, disjoint, and non-adjacent parts $S_{1}$, $S_{2}$ and satisfying $|S|\leqslant k$. Therefore, $|S_{1}|=|S_{2}|$. Hence $|V|-|S|=|S_{1}|+|S_{2}|$, or $|S|=n-2|S_{1}|$, an even integer: say $|S|=2\beta$, for some integer $\beta$. Therefore, $2\beta\leqslant 2\lambda+1\implies\beta\leqslant\lambda+0.5$. Since $\beta,\lambda$ are integers, we must have $\beta\leqslant\lambda$; implying $|S|=2\beta\leqslant 2\lambda=2\left\lfloor\frac{k}{2}\right\rfloor$. Hence $(G,2\left\lfloor\frac{k}{2}\right\rfloor)$ is also a Yes-instance.
@@
Reverse direction. We first show if the Optimal House Allocation of Agent Network instance $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}},2\left\lfloor\frac{k}{2}\right\rfloor)$ is a Yes-instance, then so is the $1/2$-Vertex Separator instance of $(G,k)$. Let $\phi$ be an allocation with at most $2\left\lfloor\frac{k}{2}\right\rfloor$ envious agents. Let $X\subseteq\mathcal{A}$ defined by $X=\{a\in\mathcal{A}\mid\phi(a)\in\mathcal{H}^{*}\}$. Therefore, $|X|=t=\frac{n}{2}-\left\lfloor\frac{k}{2}\right\rfloor$. Let $J\subseteq\mathcal{A}$ be the set of envious agents. By assumption, $|J|\leqslant 2\left\lfloor\frac{k}{2}\right\rfloor$. Let $Y=\mathcal{A}\setminus(X\cup J)$, be the set of non-envious agents outside $X$.
@@ 4.1 Single Exponential Exact Algorithms
    If $a\notin C$, then $\phi(a)\notin\mathcal{P}_{a}$. Thus $\mathcal{F}_{a}$ cannot contain any house in $\mathcal{P}_{a}$. While, for $\in C, a\mathcal{F}_{a}$ can only contain houses in $\mathcal{P}_{a}$.
@@
[…]
@@
23:   $E_{\mathcal{F}}\leftarrow\{\{a,h\}\mid a\in\mathcal{A},h\in\mathcal{F}_{a}$}; $\mathcal{G}_{\mathcal{F}}\leftarrow(\mathcal{H}\cup\mathcal{A},E_{\mathcal{F}})$ $\triangleright$ Feasibility bipartite graph